if vector a+b vector is equal to R vector and vector a+2b vector is perpendicular to a vector then 1) b=2R2) b=R

Grade:11

1 Answers

Arun
6 years ago

Here we are given two lines a+ 2b and a which are inclined at 90 degrees.

Thus applying dot product on these lines,

a× (a+2b)=0

a^2+2ab=0

a^2=-2ab

a=-2b

Now a+b=r

-2b+b=r

Thus b+r=0. Or b=-r